OpenStudy (babygirl4556):

which of the following numbers are equivalent to 20%, choose all that apply 1) 0.002 2) 1/4 3) 0.20 4) 1/5

OpenStudy (anonymous):

c and d i think but not sure

OpenStudy (er.mohd.amir):

20%=20/100 =2/10 =1/5

OpenStudy (er.mohd.amir):

=0.2